# Theory-agnostic searches for non-gravitational modes in black hole ringdown

### v2

@@ Abstract
Interestingly, despite Fig. 1 showing that there is no statistical evidence for an extra scalar or vector mode, its inclusion affects the posterior distributions of the parameters. This is shown in Table 1 for a representative example of GW150914 analyzed with an extra scalar mode. Other events, different time offsets, or the vector case show qualitatively similar results. While the presence of an extra mode does not affect the distribution of the remnant mass significantly, it contributes to broaden up that of the spin towards smaller values (see Appendix for the posterior distributions). This generic feature can be understood from the fact that an extra $(220)$ scalar/vector mode has a damping time comparable to the fundamental gravitational mode (and the damping time is only mildly sensitive to the spin for $\chi_{f}\lesssim 0.8$) and also has a higher frequency than the $(220)$ and $(221)$ gravitational modes. Thus, interpreting the overtone frequency with an extra scalar/vector mode requires a smaller remnant spin. Also, the amplitude of the fundamental gravitational mode is affected by the extra mode: in the GR0+S and GR1+S models the peak of the $A_{220}$ distribution is smaller because part of the information is contained in the scalar mode. Consequently, the amplitude ratio $A_{R,220}$ between the $(220)$ scalar mode and the ($220$) gravitational mode peaks at some nonzero value.
@@
Forecasts with future observations. In Fig. 2, we forecast the constraining power of this test by computing the minimum ringdown SNR, $\rho_{\rm det}$, for detectability of an extra scalar mode in the GR0+S model, for different amplitude ratios, remnant spins, and phase differences $\delta\phi=\phi_{220}-\hat{\phi}_{220}^{s=0}$. (Hereafter we focus on the scalar case, since the vector case gives qualitatively similar results, see Appendix.) $\rho_{\rm det}$ is defined following Refs. [85, 87, 88], namely as the SNR such that $\sigma_{A_{R}}=A_{R,220}$, where the statistical error $\sigma_{A_{R}}$ has been computed with a fully numerical Fisher information matrix [87, 88]. Note that, in contrast to overtones [70, 85], resolving the scalar mode is relatively easy since the frequency is significantly different from the gravitational one. For this reason, $\rho_{\rm det}$ is always larger than the SNR threshold required to resolve the extra scalar mode. We perform two variants of this analysis, either averaging over or fixing the sky-location of the signal, with results shown in the left and right panel of Fig. 2, respectively.
@@
$\rho_{\rm det}\approx 9.3\frac{g(\chi_{f},\delta\phi,A_{R,220})}{A_{R,220}}\,.$ (7)
where $g\approx 1+0.56A_{R,220}-0.32\chi_{f}+0.14\cos(\delta\phi+0.98)$, and for $A_{R,220}\leq 0.12$ and $\chi_{f}\leq 0.9$ the fit is accurate within $10\%$.
In the left panel of Fig. 2 we also provide reference values of the ringdown SNR for current and future detectors. In particular, third-generation ground-based detectors [89] such as Cosmic Explorer [90, 91, 92] and the Einstein Telescope [93, 94, 95] are expected to observe a few events per year with ringdown SNR greater than 100 [88]. Likewise, GW space interferometers such as LISA [96] are expected to detect up to a dozen of massive BH mergers with ringdown SNR greater than 1000, depending on the massive BH population [97, 87]. Our results show that a ringdown SNR of 150 (resp., 1000) would yield a constraint $A_{R,220}\lesssim 0.05$ (resp., 0.008) for $\chi_{f}\approx 0.9$, with only mild dependence on $\delta\phi$. Very similar results apply to the detectability of an extra vector mode.
@@
Discussion. Any theory predicting extra ringdown modes should presumably also predict deviations from the standard gravitational Kerr QNMs, in which case one could argue that GW detectors are more sensitive to phase differences (and hence to QNM shifts) rather than amplitude differences, so that our test could have less constraining power than ordinary ringdown tests. However, there are known examples of theories predicting zero or negligible QNM shifts but extra modes, in which case our test can be superior to ordinary BH spectroscopy. An example is dynamical Chern-Simons gravity, wherein for a Schwarzschild BH the GR QNMs are unchanged, but the ordinary ringdown contains also extra scalar modes [32]. In the Kerr case also the GR QNMs are modified, but the deviations are suppressed by powers of the BH spin [45], so the convenience of our method will likely depend on the remnant’s spin.
@@
As future extensions, it would be interesting to consider a specific theory and compare the constraints on the coupling constant(s) placed by our test with those of ordinary BH spectroscopy. This would require estimates of both ordinary QNM shifts and excitation amplitudes of extra modes in a given theory, both of which have recently become available for some theories [47, 48, 49, 98, 99, 100, 101, 102, 103, 104, 105]. Another possible avenue of exploration is to use the recently introduced QNM filtering technique [76, 36, 35] to search for extra (scalar, vector, etc) modes.
@@
Acknowledgements. We thank Collin Capano and Alexander Nitz for clarifications on PyCBC inference, and Gregorio Carullo, Cecilia Chirenti, Gabriele Franciolini, Nico Yunes, and the participants of the workshop Ringdown Inside and Out for insightful comments and discussions. This work is partially supported by the MUR PRIN Grant 2020KR4KN2 “String Theory as a bridge between Gauge Theories and Quantum Gravity”, by the FARE programme (GW-NEXT, CUP: B84I20000100001), and by the INFN TEONGRAV initiative. X. Jimenez is supported by the Spanish Ministerio de Ciencia, Innovación y Universidades (Beatriz Galindo, BG22-00034) and cofinanced by UIB; the Spanish Agencia Estatal de Investigación Grants No. PID2022-138626NB-I00, No. RED2022- 134204-E, and No. RED2022-134411-T, funded by MCIN/AEI/10.13039/501100011033/FEDER, UE; the MCIN with funding from the European Union NextGenerationEU/PRTR (No. PRTR-C17.I1); the Comunitat Autonòma de les Illes Balears through the Direcció General de Recerca, Innovació I Transformació Digital with funds from the Tourist Stay Tax Law (No. PDR2020/11 - ITS2017-006), and the Conselleria d’Economia, Hisenda i Innovació Grant No. SINCO2022/6719. Some numerical computations have been performed at the Vera cluster supported by the Italian Ministry of Research and by Sapienza University of Rome, and with the University of Birmingham’s BlueBEAR HPC service.
Appendix A Supplemental material
A.1 Explicit example: Dynamical Chern-Simons gravity
@@
We provide here a specific example of a nonminimal coupling giving rise to extra scalar modes in the gravitational sector. We consider a theory with quadratic curvature corrections, dynamical Chern-Simons gravity, described by the action [26]
@@
As in the GR case, the only stationary, spherically-symmetric solution is the Schwarzschild metric. For simplicity we consider perturbations of this solution, neglecting the spin of the background. Axial perturbations of the metric are coupled to those of the scalar field. Upon a spherical harmonic decomposition and in the frequency domain, they reduce to the following set of coupled ordinary differential equations [32]
$\frac{d^{2}}{dr_{\star}^{2}}\Psi+\left\{\omega^{2}-f\left[\frac{l(l+1)}{r^{2}}-\frac{6M}{r^{3}}\right]\right\}\Psi$ $=$ $\frac{96\pi\,Mf}{r^{5}}\alpha\Theta\,,$ (9)
$\frac{d^{2}}{dr_{\star}^{2}}\Theta+\left\{\omega^{2}-f\left[\frac{l(l+1)}{r^{2}}\left(1+\frac{576\pi M^{2}\alpha^{2}}{r^{6}}\right)+\frac{2M}{r^{3}}\right]\right\}\Theta$ $=$ $f\frac{(l+2)!}{(l-2)!}\frac{6M\alpha}{r^{5}}\Psi$ (10)
where $f(r)=1-2M/r$ and $r_{\star}\equiv r+2M\ln\left(r/2M-1\right)$ is the standard Schwarzschild tortoise coordinate. The variables $\Psi$ and $\Theta$ reduce to the standard metric and scalar master functions, respectively, in the decoupling limit, $\alpha\to 0$. Indeed, when $\alpha=0$ the two equations decouple and reduce to the standard Regge-Wheeler equation and scalar-perturbation equation of a Schwarzschild BH, respectively. However, when $\alpha\neq 0$ the two perturbations are coupled to each other and the scalar effective potential acquires some corrections.
The coupling $\alpha$ gives rise to two features [32, 34]:
@@
    The above system of equations is akin to a coupled harmonic oscillator, so a scalar perturbation would source scalar modes in the gravitational sector, and vice versa. In particular, in this theory the scalar field is at least linear in $\alpha$ [26], in which case the coupled perturbation equations imply that, whether or not scalar perturbations are present in the merger conditions (for example if the progenitors are endowed with a scalar field [106]), the amplitude of the scalar mode in the gravitational sector would be ${\cal O}(\alpha^{2})$. In the notation of the main text, this would imply that, at the leading order, $A_{R,220}\propto\alpha^{2}/M_{f}^{4}$, which is the same order of the usual leading-order corrections to the QNMs in this theory [32, 45, 49].
@@
where the sum runs over the overtones. The scalar modes $\hat{\omega}_{j}$ contains ${\cal O}(\alpha^{2})$ corrections, but $\hat{A}_{j}$ is at least ${\cal O}(\alpha)$ or higher. Therefore, to leading order in $\alpha$ we can approximate $\hat{\omega}_{j}\approx\omega_{j}^{\rm GR,~{}scal}$ in the second term of the above equation, which is the crucial simplification of our test. This is consistent with the finding of Ref. [32], where the gravitational ringdown in the $\alpha\ll M^{2}$ limit contains both the unperturbed gravitational and scalar modes.
While we explicitly showed this for a specific theory, it is in fact a very general properties of extended theories of gravity with nonminimal couplings (see, e.g., [56] for another example).
A.2 The effect of precession on GW190521 and GW200129

### v3

@@ III Bayesian analysis on real data
Interestingly, despite Fig. 1 showing that there is no statistical evidence for an extra scalar or vector mode, its inclusion affects the posterior distributions of the parameters. This is shown in Table 1 for a representative example of GW150914 analyzed with an extra scalar mode, where we denote the scalar-to-tensor mode amplitude ratio as $A_{R,220}=\hat{A}_{220}^{s=0}/A_{220}$. These results are also consistent with the full inspiral-merger-ringdown constraints provided in [LIGOScientific:2021sio]. However, as expected, our peak values are shifted compared to those computed using a GR waveform approximant. Other events, different time offsets, or the vector case show qualitatively similar results. While the presence of an extra mode does not affect the distribution of the remnant mass significantly, it contributes to broadening that of the spin towards smaller values (see Appendix VII.3 for the posterior distributions). This generic feature can be understood from the fact that an extra $(220)$ scalar/vector mode has a damping time comparable to the fundamental gravitational mode (and the damping time is only mildly sensitive to the spin for $\chi_{f}\lesssim 0.8$) and also has a higher frequency than the $(220)$ and $(221)$ gravitational modes. Thus, interpreting the overtone frequency with an extra scalar/vector mode requires a smaller remnant spin. Also, the amplitude of the fundamental gravitational mode is affected by the extra mode: in the GR0+S and GR1+S models the peak of the $A_{220}$ distribution is smaller because part of the information is contained in the scalar mode. Consequently, the amplitude ratio $A_{R,220}$ between the $(220)$ scalar mode and the ($220$) gravitational mode peaks at some nonzero value. Finally, we do not find strong support for the presence of an additional mode for $t_{\text{offset}}>-2\mathrm{ms}$. As in previous related work [Cotesta:2022pci, Isi:2023nif], we observe that the Bayes factor tends to increase when the analysis is extended to negative times, since the model begins to fit higher overtones and nonlinear contributions. We interpret this increase not as evidence for extra modes, but rather as a consequence of model mismatch arising from applying the ringdown model before the merger, which can yield spurious results. Furthermore, for the event GW200129 – where we observe the largest values of $\log_{10}\mathcal{B}^{A}_{\text{GR1}}$ – data quality issues [Payne:2022spz] and possible signatures of orbital eccentricity and/or precession [Hannam:2021pit, Gupte:2024jfe] are known to affect the inference of physical parameters. These factors could also contribute to the early-time rise of the Bayes factor in support for additional modes. A detailed investigation of these effects will be presented in future work.
IV Forecasts with future observations
@@
In Fig. 2, we forecast the constraining power of this test by computing the minimum ringdown SNR, $\rho_{\rm det}$, for detectability of an extra scalar mode in the GR0+S model, for different amplitude ratios, remnant spins, and phase differences $\delta\phi=\phi_{220}-\hat{\phi}_{220}^{s=0}$. (Hereafter we focus on the scalar case, since the vector case gives qualitatively similar results, see Appendix VII.3.) $\rho_{\rm det}$ is defined following Refs. [JimenezForteza:2020cve, Bhagwat:2021kwv, Bhagwat:2023jwv], namely as the SNR such that $\sigma_{A_{R}}=A_{R,220}$, where the statistical error $\sigma_{A_{R}}$ has been computed with a fully numerical Fisher information matrix , also assuming that the relevant sensitivity curves at the ringdown regime are approximately flat for the two distinct final masses considered [Bhagwat:2021kwv, Bhagwat:2023jwv]. In general, beyond-GR theories can introduce up to six polarization modes, by adding extra antenna-pattern response terms to the signal $h(t)$ [Yunes:2024lzm]. Instead, we focus here on the corrections to the standard $_{\times}, h+h_{+}$ GR polarizations, which are already detectable with current GW networks. Note that, in contrast to overtones [Bhagwat:2019dtm, JimenezForteza:2020cve], resolving the frequency of the scalar mode is relatively easy since it is significantly different from the gravitational one. For this reason, $\rho_{\rm det}$ is always larger than the SNR threshold required to resolve the extra scalar mode. The requirements for the detectability and resolvability of a mode are outlined in the Appendix VII.4.
We perform two variants of this analysis, either averaging over or fixing the sky-location of the signal, with results shown in the left and right panel of Fig. 2, respectively. Notice that, since the SNR is fully correlated with the sky location, both approaches should yield consistent results (see Appendix VII.4). Therefore, the right panel simply represents a phase-expanded version of the left-hand side plot.
@@
$\rho_{\rm det}\approx{4.0}\frac{g(\chi_{f},\delta\phi,A_{R,220})}{A_{R,220}}\,.$ (7)
where $g\approx 1+0.25A_{R,220}-0.40\chi_{f}+0.08\cos(\delta\phi+1.16)$, and for $A_{R,220}\leq 0.12$ and $\chi_{f}\leq 0.9$ the fit is accurate within $10\%$.
In the left panel of Fig. 2 we also provide reference values of the ringdown SNR for current and future detectors. In particular, third-generation ground-based detectors [Kalogera:2021bya] such as Cosmic Explorer [Reitze:2019iox, Evans:2021gyd, Evans:2023euw] and the Einstein Telescope [Hild:2010id, Maggiore:2019uih, Branchesi:2023mws] are expected to observe a few events per year with ringdown SNR greater than 100 [Bhagwat:2023jwv]. Likewise, GW space interferometers such as LISA [Colpi:2024xhw] are expected to detect up to a dozen of massive BH mergers with ringdown SNR greater than 1000, depending on the massive BH population [Berti:2016lat, Bhagwat:2021kwv]. Our results show that a ringdown SNR of 150 (resp., 1000) would yield a constraint $A_{R,220}\lesssim 0.02$ (resp., 0.003) for $\chi_{f}\approx 0.9$, with only mild dependence on $\delta\phi$. Very similar results apply to the detectability of an extra vector mode.
@@ V Discussion
Any theory predicting extra ringdown modes should presumably also predict deviations from the standard gravitational Kerr QNMs, in which case one could argue that GW detectors are more sensitive to phase differences (and hence to QNM shifts) rather than amplitude differences, so that our test could have less constraining power than ordinary ringdown tests. However, there are known examples of theories predicting zero or negligible QNM shifts but extra modes, in which case our test can be superior to ordinary BH spectroscopy. An example is dynamical Chern-Simons gravity, wherein for a Schwarzschild BH the polar GR QNMs are unchanged, but the ordinary ringdown contains also extra scalar modes [Molina:2010fb]. In the Kerr case also all the GR QNMs are modified, but the deviations are suppressed by powers of the BH spin [Wagle:2021tam], so the convenience of our method will likely depend on the remnant’s spin.
The parametrization introduced in this work can also apply to cases in which the remnant is a Kerr BH or in the presence of extra fields but still within the realm of GR, provided the deviations from the Kerr case are sufficiently small. For example, it applies to Kerr–Newman BHs with small charge $Q_{f}$. In the perturbative limit, $Q_{f}/M_{f}\ll 1$, the parameter $Q_{f}/M_{f}$ effectively plays the role of a coupling constant in an effective field theory, making Eq. (I) applicable in this scenario, upon replacing the scalar sector with that of a test spin-$1$ field.
We focused on $\ell=m=2$ but clearly, depending on the source, one should also add $\ell=3$ modes, or overtones. In this context, it is noteworthy that the ringdown analysis for the latest GW event, GW250114 [LIGOScientific:2025obp, LIGOScientific:2025rid], was mainly performed precisely with $\ell=m=2$ modes only, using the fundamental mode and the first overtone. We point out that, for events such as GW250114 (but also for GW150914) a ringdown tests of GR can be based on $\ell=m=2$ only, but it would be inconsistent to include QNM shifts alone, without adding extra modes. Also, note that nearly equal-mass binaries are expected to dominate ringdown tests in both current and future GW detectors. For instance, using the fits of Ref. [JimenezForteza:2020cve], binaries that are nearly face-on ($\iota<30^{\circ}$) with mass ratios $q\lesssim 1.2$ and effective spins $\chi_{\rm eff}\sim 0.7$ exhibit amplitude ratios of approximately $(A_{33}Y_{33}/A_{22}Y_{22})\sim 0.03$ and $(A_{22,{\rm s}}Y_{22}/A_{22}Y_{22})=A_{22,{\rm s}}/A_{22}$. Therefore, values of $A_{22,{\rm s}}/A_{22}$ as low as $\sim 0.03$ would still produce a secondary component comparable in strength to the $(3,3)$ mode. A similar argument applies to the $(2,1)$ mode, since in the same parameter regime one finds $(A_{21}Y_{21}/A_{22}Y_{22})\sim 0.03$. Regarding overtones, given their short damping times, their relevance can be reduced by starting the ringdown analysis later (when scalar modes are still relevant). Clearly, this would reduce the overall SNR of the effective ringdown used for the test but would also avoid the data-analysis subtleties associated with overtones [Cotesta:2022pci, Baibhav:2023clw].
@@
As future extensions, it would be interesting to consider a specific theory and compare the constraints on the coupling constant(s) placed by our test with those of ordinary BH spectroscopy. This would require estimates of both ordinary QNM shifts and excitation amplitudes of extra modes in a given theory, both of which have recently become available for some theories [Pierini:2022eim, Cano:2023jbk, Wagle:2023fwl, Okounkova:2019zjf, Okounkova:2020rqw, East:2020hgw, East:2021bqk, Figueras:2021abd, AresteSalo:2022hua, Corman:2022xqg, Cayuso:2023aht]. As an order-of-magnitude estimate, for a theory which adds quadratic curvature corrections to the standard Einstein-Hilbert action, the coupling constant $\alpha$ has the dimension of a length squared [Berti:2015itd] In such a case one would expect $A_{R,220}={\cal O}(1)\alpha^{2}/M_{f}^{4}$. Our results suggest that future detectors will be able to probe $A_{R,220}={\cal O}(10^{-3})$, and hence a coupling constant as small as $\alpha^{2}/M_{f}^{4}\sim{\cal O}(10^{-3})$.
A related extension is to include higher harmonics in the test. Besides adding complexity to the mode, this would also require including the inclination and the sky localization as extra waveform parameters. Work in this direction is underway.
Finally, another possible avenue of exploration is to use the recently introduced QNM filtering technique [Ma:2023vvr, Ma:2022wpv, Ma:2023cwe] to search for extra (scalar, vector, etc) modes.
VI Acknowledgements
@@
We thank Collin Capano and Alexander Nitz for clarifications on PyCBC inference, and Gregorio Carullo, Cecilia Chirenti, Gabriele Franciolini, Nicolas Yunes, and the participants of the workshop Ringdown Inside and Out for insightful comments and discussions. This work is partially supported by the MUR PRIN Grant 2020KR4KN2 “String Theory as a bridge between Gauge Theories and Quantum Gravity”, by the FARE programme (GW-NEXT, CUP: B84I20000100001), and by the INFN TEONGRAV initiative. X. Jimenez is supported by the Spanish Ministerio de Ciencia, Innovación y Universidades (Beatriz Galindo, BG22-00034) and cofinanced by UIB; the Spanish Agencia Estatal de Investigación Grants No. PID2022-138626NB-I00, No. RED2022- 134204-E, and No. RED2022-134411-T, funded by MCIN/AEI/10.13039/501100011033/FEDER, UE; the MCIN with funding from the European Union NextGenerationEU/PRTR (No. PRTR-C17.I1); the Comunitat Autonòma de les Illes Balears through the Direcció General de Recerca, Innovació I Transformació Digital with funds from the Tourist Stay Tax Law (No. PDR2020/11 - ITS2017-006), and the Conselleria d’Economia, Hisenda i Innovació Grant No. SINCO2022/6719. Some numerical computations have been performed at the Vera cluster supported by the Italian Ministry of Research and by Sapienza University of Rome, and with the University of Birmingham’s BlueBEAR HPC service.
VII Appendix
VII.1 Explicit example: Dynamical Chern-Simons gravity
@@
We provide here a specific example of a nonminimal coupling giving rise to extra scalar modes in the gravitational sector. We consider a theory with quadratic curvature corrections, dynamical Chern-Simons gravity, described by the action [Alexander:2009tp]
@@
As in the GR case, the only stationary, spherically-symmetric solution is the Schwarzschild metric. For simplicity we consider perturbations of this solution, neglecting the spin of the background. Axial perturbations of the metric are coupled to those of the scalar field. Upon a spherical harmonic decomposition and in the frequency domain, they reduce to the following set of coupled ordinary differential equations [Molina:2010fb]
$\begin{cases}\frac{d^{2}}{dr_{\star}^{2}}\Psi+\left\{\omega^{2}-f\left[\frac{l(l+1)}{r^{2}}-\frac{6M}{r^{3}}\right]\right\}\Psi=\frac{96\pi Mf}{r^{5}}\alpha\Theta,\\[10.0pt] \frac{d^{2}}{dr_{\star}^{2}}\Theta+\left\{\omega^{2}-f\left[\frac{l(l+1)}{r^{2}}\left(1+\frac{576\pi M^{2}\alpha^{2}}{r^{6}}\right)+\frac{2M}{r^{3}}\right]\right\}\Theta=f\frac{(l+2)!}{(l-2)!}\frac{6M\alpha}{r^{5}}\Psi\end{cases}$ (9)
where f(r)=12M/rf(r)=1-2M/r and $r_{\star}\equiv r+2M\ln\left(r/2M-1\right)$ is the standard Schwarzschild tortoise coordinate. The variables $\Psi$ and $\Theta$ reduce to the standard metric and scalar master functions, respectively, in the decoupling limit, $\alpha\to 0$. Indeed, when $\alpha=0$ the two equations decouple and reduce to the standard Regge-Wheeler equation and scalar-perturbation equation of a Schwarzschild BH, respectively. However, when $\alpha\neq 0$ the two perturbations are coupled to each other and the scalar effective potential acquires some corrections.
The coupling $\alpha$ gives rise to two features [Molina:2010fb, Cardoso:2020nst]:
@@
    The above system of equations is akin to a coupled harmonic oscillator, so a scalar perturbation would source scalar modes in the gravitational sector, and vice versa. In particular, in this theory the scalar field is at least linear in $\alpha$ [Alexander:2009tp], in which case the coupled perturbation equations imply that, whether or not scalar perturbations are present in the merger conditions (for example if the progenitors are endowed with a scalar field [Yunes:2009hc]), the amplitude of the scalar mode in the gravitational sector would be ${\cal O}(\alpha^{2})$. In the notation of the main text, this would imply that, at the leading order, $A_{R,220}=\gamma\cdot\alpha^{2}/M_{f}^{4}$, where $\gamma$ is a source-dependent excitation factor [Molina:2010fb, Wagle:2021tam, Wagle:2023fwl, Crescimbeni:2025kxi].
@@
where the sum runs over the overtones. The scalar modes $\hat{\omega}_{j}$ contains ${\cal O}(\alpha^{2})$ corrections, but $\hat{A}_{j}$ is at least ${\cal O}(\alpha)$ or higher. Therefore, to leading order in $\alpha$ we can approximate $\hat{\omega}_{j}\approx\omega_{j}^{\rm GR,~scal}$ in the second term of the above equation, which is the crucial simplification of our test. This is consistent with the finding of Ref. [Molina:2010fb], where the gravitational ringdown in the $\alpha\ll M^{2}$ limit contains both the unperturbed gravitational and scalar modes.
While we explicitly showed this for a specific theory, it is in fact a very general properties of extended theories of gravity with nonminimal couplings (see, e.g., [Blazquez-Salcedo:2016enn] for another example).
VII.2 The effect of precession on GW190521 and GW200129